A truck slow down it's speed from 216km/h to 18m/h in 0.25minutes

Calculate the truck acceleration

Fine the distance over in meter


1
Expert's answer
2021-03-14T19:21:26-0400

The truck acceleration can be calculated as follows:


"a=\dfrac{v-v_0}{t},""a=\dfrac{18\ \dfrac{m}{s}-216\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s}}{0.25\cdot60\ s}=-2.8\ \dfrac{m}{s^2}."

The sign minus means that the truck decelerates.

The distance traveled by the truck can be found as follows:


"d=\dfrac{1}{2}(v+v_0)t,""d=\dfrac{1}{2}\cdot(18\ \dfrac{m}{s}+60\ \dfrac{m}{s})\cdot15\ s=585\ m."